# Qingdao 266237, China

### v1

@@ Appendix 0.E Missing Security Proofs
    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$. Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For the view of $P_{c}(c\neq 1,c\neq m)$, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$’s view, it runs the protocol honestly with the following changes:
@@
    • It simulates uniform outputs $r_{i,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniform messages $u_{i,m}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}=r_{m}$, where $P_{i}\in\mathcal{H}$.
@@
    Clearly, the joint distribution of $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
0.E.2 The Proof of Theorem 5.2
@@
For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$, $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$ as its share, and $m-1$ sets of shares $\{\vec{sh}_{i}^{\prime}\}_{1<i\leq m}$($P_{i}$’s output from $\mathcal{F}_{\mathsf{ms}}$) from $P_{i}$ if $c=1$. The simulator emulates each $P_{c}$’s view by running the protocol honestly with the following changes:
@@
  • In step 3, it simulates uniform outputs $\{r_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j},\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, on condition that $\exists 1\leq d\leq j,P_{d}\in\mathcal{H}$.
@@
The changes of outputs from $\mathcal{F}_{\mathsf{bssPMT}}$ and $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ have no impact on $P_{c}$’s view, for the following reasons. By the definition of $\mathcal{F}_{\mathsf{bssPMT}}$, each output $e_{c,j}^{b}$ and $e_{c,i}^{b}$ from $\mathcal{F}_{\mathsf{bssPMT}}$ is uniformly distributed as a secret-share between $P_{c}$ and $P_{j}$, or $P_{i}$ and $P_{c}$, where $P_{i},P_{j}\in\mathcal{H}$. By the definition of $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, each output $r_{c,ji}^{b}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ is a secret-share of 00 among $P_{1},\cdots,P_{j}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=0$, or a secret-share of $\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. Therefore, even if $P_{c}$ colludes with others, $r_{c,ji}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one share. Besides, the outputs $\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ are uniformly distributed.
@@
We now proceed to explain the latter case. For all $1<j\leq m$, consider an element $x\in X_{j}$ and $x$ is placed in the $b$th bin by $P_{j}$. In the real protocol, if there is no $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then for all $1\leq i<j$, $e_{i,j}^{y}\oplus e_{j,i}^{y}=0$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=0$. From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)$, namely, they are additive shares of $x\|\mathsf{H}(x)$ among all parties. This is exactly identical to the simulation.
If there exists some $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$, where each $\Delta_{d,ji}^{b}$ is uniformly held by $P_{d}$. From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)\oplus\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}\oplus r$, where $r$ is the sum of remaining terms. Then, even if $P_{1}$ colludes with others, $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one uniform $\Delta_{d,ji}^{b}$ and independent of all honest parties’ inputs. For all empty bins, $u_{d,j}^{b}$ is chosen uniformly random, so the corresponding $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is also uniformly at random, which is identical to the simulation. By the definition of $\mathcal{F}_{\mathsf{ms}}$, all parties additively share $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ in a random permutation that maintains privacy against a coalition of arbitrary corrupted parties, and receive back $\{\vec{sh}_{i}^{\prime}\}$, respectively. We conclude that all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$ distribute identically between the real and ideal executions.
0.E.3 The Proof of Theorem 6.1
@@
Appendix 0.F Complete Complexity Analyses of Protocols
0.F.1 Theoretical Analysis of Protocol $\Pi_{\mathsf{bssPMT}}$
@@
      For $B=O(n)$ instances of OPRF, we require performing a subfield-VOLE of size $B$ in the offline phase, and sending $B$ derandomization messages of length $l$ in the online phase. We resort to Dual LPN with Fixed Weight and Regular Noise [BCG+19a, BCG+19b] to improve the efficiency of subfield-VOLE, and instantiate the code family with Expand-Convolute codes presented by [RRT23], which enables near-linear time syndrome computation. The computation complexity is $O(n\log n)$. The computation complexity is $O(t\lambda\log n/t)$, where $t$ is the fixed noise weight [^20] For instance, $t\approx 128$.. The round complexity is $O(1)$.
    • OKVS: The size of key-value pairs encoded into OKVS is $\lvert K_{1}\rvert+\cdots+\lvert K_{B}\rvert$. Taking into account the invocation of batch ssPMT in our two MPSU protocols in advance, this size is $3n$. We use the OKVS construction of [RR22], and the computation complexities of $\mathsf{Encode_{H}}$ and $\mathsf{Decode_{H}}$ algorithms are both $O(n)$. As we employ their $w=3$ scheme with a cluster size of $2^{14}$, the size of OKVS is $1.28\gamma\cdot(3n)$ bits, where $\gamma$ is the output length of OPPRF and is set as $\sigma+\log_{2}B$ to avoid wrong matching.
  • Offline phase. The computation complexity of each party is $O(n\log n)$. The communication complexity of each party is $O(t\lambda\log(n/t))$. The round complexity is $O(1)$.
  • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(1)$.
@@
    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
  • 0.F.2 Theoretical Analysis of Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$
@@
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
  • Multi-party secret-shared random oblivious transfers. Each invocation of mss-ROT involves pairwise executions of two-party OT. If $P_{i}$ and $P_{j}$ hold the two choice bits, then $P_{i}$ and $P_{j}$ invoke two instances of OT and they separately invoke one instance of OT with the remaining parties.
    Each OT execution consists of two parts: In the offline phase, the parties engages in a random-choice-bit ROT, then $\mathcal{S}$ sends one messages of length $l+\kappa$ ($\kappa=\sigma+\log(m-1)+\log n$) to $\mathcal{R}$ [^22] The OT execution between two choice-bit-holder does not include the sending step. This does not affect the total complexity, so we ignore this.; In the online phase, $\mathcal{R}$ sends a 1-bit derandomization message to transform ROT into a chosen-input version.
    For $1\leq i<j\leq m$, $P_{1},\cdots,P_{j}$ engage in $B$ instances of mss-ROT, where $P_{i}$ and $P_{j}$ hold the two choice bits. Considering one instance of mss-ROT and fixing $j$, there are overall $j-1$ invocations of mss-ROT. For $1\leq d<j$, each $P_{d}$ has one chance to act as the choice-bit-holder (who invokes two instances of OT with another and one with $j-2$ remaining parties separately in one invocation of mss-ROT) and acts as a nomal party (who invokes one instances of OT with the two choice-bit-holders separately in one invocation of mss-ROT) in the rest $j-2$ invocations. $P_{j}$ act as the choice-bit-holder in $j-1$ invocations (who invokes $2+(j-2)=j$ instances of OT).
    Overall, each party $P_{j^{\prime}}$ invokes $[{j^{\prime}}({j^{\prime}}-1)+\sum_{j={j^{\prime}}+1}^{m}(j+2(j-2))]B=(\frac{3m^{2}-{j^{\prime}}^{2}-5m+3{j^{\prime}}}{2})B$ instances of two-party OT.
    • Offline phase. The computation complexity of each party is $O(m^{2}n\log n)$. The communication complexity of each party is $O(t\lambda m^{2}\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(m^{2}n)$. The communication complexity of each party is $O(m^{2}n)$. The round complexity is $O(1)$.
@@
    • Offline phase. The computation complexity of each party is $O(m^{2}n\log(mn))$. The communication complexity of each party is $O(\lambda m^{2}n\log(mn))$. The round complexity is $O(1)$.
@@
  • 0.F.3 Theoretical Analysis of Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$

### v4

@@ Appendix 0.D Missing Security Proofs
      For $P_{1}$’s view, the simulator runs the protocol honestly except that it simulates uniform outputs $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For the view of $P_{c}(c\neq 1,c\neq m)$, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$’s view, it runs the protocol honestly with the following changes:
@@
      • It simulates uniform outputs $r_{i,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniform messages $u_{i,m}$ from $P_{i}$ under the constraint $\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m}=r_{m}$, where $P_{i}\in\mathcal{H}$.
@@
      Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
    0.D.2 The Proof of Theorem 5.2
@@
    For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$, $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$ as its share, sampled values as shares of $\Delta$ for $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ and $m-1$ sets of shares $\{\vec{sh}_{i}^{\prime}\}_{1<i\leq m}$($P_{i}$’s output from $\mathcal{F}_{\mathsf{ms}}$) from $P_{i}$ if $c=1$. The simulator emulates each $P_{c}$’s view by running the protocol honestly with the following changes:
@@
    • In step 3, it simulates uniform outputs $\{r_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, on condition that $\exists min(2,i)\leq d\leq j,P_{d}\in\mathcal{H}$. If $c\neq 1$, it simulates uniform $\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ as $P_{c}$’s random tapes.
@@
    The changes of outputs from $\mathcal{F}_{\mathsf{bssPMT}}$ and $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ have no impact on $P_{c}$’s view, for the following reasons. By the definition of $\mathcal{F}_{\mathsf{bssPMT}}$, each output $e_{c,j}^{b}$ and $e_{c,i}^{b}$ from $\mathcal{F}_{\mathsf{bssPMT}}$ is uniformly distributed as a secret-share between $P_{c}$ and $P_{j}$, or $P_{i}$ and $P_{c}$, where $P_{i},P_{j}\in\mathcal{H}$. By the definition of $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, each output $r_{c,ji}^{b}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ is a secret-share of 00 among $P_{min(2,i)},\cdots,P_{j}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=0$, or a secret-share of $\bigoplus_{d=2}^{j}\Delta_{d,ji}^{b}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. Therefore, even if $P_{c}$ colludes with others, $r_{c,ji}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(min(2,i)\leq d\leq j)$ holding one share.
@@
    We now proceed to explain the latter case. For all $1<j\leq m$, consider an element $x\in X_{j}$ and $x$ is placed in the $b$th bin by $P_{j}$. In the real protocol, if there is no $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then for all $1\leq i<j$, $e_{i,j}^{y}\oplus e_{j,i}^{y}=0$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 9, each $r_{d,ji}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=min(2,i)}^{j}r_{d,ji}^{b}=0$. From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)$, namely, they are additive shares of $x\|\mathsf{H}(x)$ among all parties. This is exactly identical to the simulation.
    If there exists some $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 9, each $r_{d,ji}^{b}$ is uniform conditioned on $\bigoplus_{d=min(2,i)}^{j}r_{d,ji}^{b}=\bigoplus_{j^{\prime}=2}^{j}\Delta_{j^{\prime},ji}^{b}$, where each $\Delta_{j^{\prime},ji}^{b}$ is uniformly held by $P_{j^{\prime}}$ ($1<j^{\prime}\leq j$). From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform conditioned on $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)\oplus\bigoplus_{j^{\prime}=2}^{j}\Delta_{j^{\prime},ji}^{b}\oplus r$, where $r$ is the sum of remaining terms. Then, even if $P_{1}$ colludes with others, $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{j^{\prime}}\in\mathcal{H}(1<j^{\prime}\leq j)$ holding one uniform $\Delta_{j^{\prime},ji}^{b}$ and independent of all honest parties’ inputs. For all empty bins, $u_{d,j}^{b}$ is chosen uniformly random, so the corresponding $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is also uniformly at random, which is identical to the simulation. By the definition of $\mathcal{F}_{\mathsf{ms}}$, all parties additively share $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ in a random permutation that maintains privacy against a coalition of arbitrary corrupted parties, and receive back $\{\vec{sh}_{i}^{\prime}\}$, respectively. We conclude that all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$ distribute identically between the real and ideal executions.
    0.D.3 The Proof of Theorem 6.1
@@
    Appendix 0.E Complexity Analyses and Comparisons
    0.E.1 Complete Analysis of LG
    The costs of each stage in LG are calculated as follows.
    • mq-ssPMT. The cost of this stage consists of three parts:
      • Offline phase. The computation complexity per party is $O((T+l-\log n)mn\log((T+l-\log n)n))$. The communication complexity per party is $O(t\lambda m\log(((T+l-\log n)n)/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity per party is $O((T+l-\log n)mn)$. The communication complexity per party is $O((T+l+\kappa-\log n)mn)$. The round complexity is $O(\log(l-\log n))$.
    • Random oblivious transfers. For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke ROT extension of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ instances of ROT extension of size $B$.
      • Offline phase. The computation complexity per party is $O(mn\log n)$. The communication complexity per party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(mn)$. The round complexity is $O(1)$.
      The costs of the remaining two steps, multi-party secret-shared shuffle and output reconstruction, are exactly the same as those in our SK-MPSU.(cf. Appendix  0.E.3).
    • Total costs.
      • Offline phase. The computation complexity per party is $O((T+l-\log n)mn\log((T+l-\log n)n)+m^{2}n\log mn)$. The communication complexity per party is $O(t\lambda m\log(((T+l-\log n)n)/t)+\lambda m^{2}n(\log m+\log n))$. The round complexity is $O(1)$.
    0.E.2 Theoretical Analysis of Batch ssPMT
    0.E.2.1 Complete Analysis
@@
        For $B=O(n)$ instances of OPRF, we require performing a subfield-VOLE of size $B$ in the offline phase, and sending $B$ derandomization messages of length $l$ in the online phase. We resort to Dual LPN with Fixed Weight and Regular Noise [BCG+19a, BCG+19b] to improve efficiency of subfield-VOLE, and instantiate the code family with Expand-Convolute codes [RRT23], which enables near-linear time syndrome computation. The computation complexity is $O(n\log n)$. The computation complexity is $O(t\lambda\log n/t)$, where $t$ is the fixed noise weight [^20] For instance, $t\approx 128$.. The round complexity is $O(1)$.
        We denote the output length of OPPRF as $\gamma$. The lower bound of $\gamma$ is relevant to the total number of the batch OPPRF invocations. In our SK-MPSU and PK-MPSU, for $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke batch OPPRF. Overall, there are $21+2+\cdots+(m-1)=(m^{2}-m)/$ 2 invocations of batch OPPRF. Considering all these invocations, we set $\gamma\geq\sigma+\log((m^{2}-m)/2)+\log_{2}B$, so that the probability of any $t_{i}\neq s_{i}$ occurring if $x\notin X_{i}$, which is $((m^{2}-m)/2))B\cdot 2^{-\gamma}$, is less than or equal to $2^{-\sigma}$.
      • OKVS: The size of key-value pairs encoded into OKVS is $\lvert K_{1}\rvert+\cdots+\lvert K_{B}\rvert$. Taking into account the invocation of batch ssPMT in our two MPSU protocols in advance, this size is $3n$. We use the OKVS construction of [RR22], and the computation complexities of $\mathsf{Encode_{H}}$ and $\mathsf{Decode_{H}}$ algorithms are both $O(n)$. As we employ their $w=3$ scheme with a cluster size of $2^{14}$, the size of OKVS is $1.28\gamma\cdot(3n)$ bits.
    • Offline phase. The computation complexity per party is $O(n\log n)$. The communication complexity per party is $O(t\lambda\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(n)$. The communication complexity per party is $O(\gamma n)$. The round complexity is $O(1)$.
@@
    • Offline phase. The computation complexity per party is $O(\gamma n\log n)$. The communication complexity per party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(n)$. The communication complexity per party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
  • 0.E.2.2 Comparison with mq-ssPMT
    Liu and Gao adapt the SKE-based mq-RPMT in [ZCL+23] to construct mq-ssPMT. The most expensive part of their construction is secret-shared oblivious decryption-then-matching (ssVODM) [ZCL+23], which is to implement a decryption circuit and a comparison circuit by the GMW protocol [GMW87]. In total, the ssVODM requires $(T+l-\log n-1)n$ AND gates, where $l$ is the bit length of set elements, and $T$ is the number of AND gates in the SKE decryption circuit and is set to be considerably large ($\approx 600$) according to their paper.
    The costs of batch ssPMT of size $B=1.27n$ [^22] We use stash-less Cuckoo hashing [PSTY19] with 3 hash functions, where $B=1.27n$. consist of the costs of batch OPPRF of size $1.27n$ and the costs of $1.27n$ instances of ssPEQT, where batch OPPRF can be instantiated by extremely fast specialized protocol and ssPEQT is implemented by the GMW protocol. Moreover, the state-of-the-art batch OPPRF construction [CGS22, RS21, RR22] can achieve linear computation and communication with respect to $n$ In ssPEQT, the parties engage in $1.27(\gamma-1)n$ AND gates, where $\gamma$ is the output length of OPPRF. In the typical setting where $n\leq 2^{24},l\leq 128,\gamma\leq 64$, we have $(T+l-\log n-1)n\gg 1.27(\gamma-1)n$, which means that the number of AND gates desired involved in batch ssPMT is far smaller than mq-ssPMT. Therefore, the construction built on batch ssPMT greatly reduces the dependency on general 2PC and significantly decreases both computation and communication complexity by a considerable factor, compared to mq-ssPMT.
    0.E.3 Theoretical Analysis of Our SK-MPSU
    0.E.3.1 Complete Analysis
@@
    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
  • Multi-party secret-shared random oblivious transfers. Each invocation of mss-ROT involves pairwise executions of two-party OT, where each OT execution consists of two parts: In the offline phase, the parties invoke random-choice-bit ROT, then $\mathcal{S}$ sends one messages of length $l+\kappa$ ($\kappa=\sigma+\log(m-1)+\log n$) to $\mathcal{R}$; In the online phase, $\mathcal{R}$ sends a 1-bit derandomization message to transform ROT into a chosen-input version. If $P_{i}$ and $P_{j}$ hold choice bits, then $P_{i}$ and $P_{j}$ invoke OT with the parties in $\{P_{2},\cdots,P_{j}\}$ except itself.
    For $1\leq i<j\leq m$, $P_{min(2,i)},\cdots,P_{j}$ engage in $B$ instances of mss-ROT, where $P_{i}$ and $P_{j}$ hold the two choice bits. Considering the overall invocations of mss-ROT: $P_{1}$ invokes $[\sum_{j=2}^{m}(j-1)]B=\frac{m^{2}B-mB}{2}$ instances of two-party OT; For $1<i\leq m$, each party $P_{i}$ invokes $[(i-1)+\sum_{j=i+1}^{m}3(j-2)]B=\frac{3m^{2}B-3i^{2}B-9mB+11iB-2B}{2}$ instances of two-party OT.
    • Offline phase. The computation complexity per party is $O(m^{2}n\log n)$. The communication complexity per party is $O(t\lambda m^{2}\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(m^{2}n)$. The communication complexity per party is $O(m^{2}n)$. The round complexity is $O(1)$.
@@
    • Offline phase. The computation complexity per party is $O(m^{2}n\log(mn))$. The communication complexity per party is $O(\lambda m^{2}n\log(mn))$. The round complexity is $O(1)$.
@@
  • 0.E.3.2 Comparison with LG
    We present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between LG and our SK-MPSU.
    Computation Communication
    Offline Online Offline Online
    Leader Client Leader Client
    Table 3: Asymptotic communication (bits) and computation costs of LG and our SK-MPSU in the offline and online phases. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $T$ is the number of AND gate in an SKE decryption circuit, $T\approx 600$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
    Computation Communication
    Offline Online Offline Online
    Ours $\gamma mn\log n$ $mn$ $t\lambda m\log n$ $(\gamma+\lambda)mn$
    Table 4: Asymptotic communication (bits) and computation costs of [GNT23] and our PK-MPSU in the offline and online phases. In the offline phase, the computation is composed of symmetric-key operations; In the online phase, the computation is composed of public-key operations since we ignore symmetric-key operations. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
    0.E.4 Theoretical Analysis of Our PK-MPSU
    0.E.4.1 Complete Analysis